OpenStudy (anonymous):

X +2. X - 2 -------- = -------- 2. 5 Slove

OpenStudy (anonymous):

\[\frac{x+2 }{ 2 }=\frac{ x -2 }{ 5 }\\\\ (10)\frac{x+2 }{ 2 }=(10)\frac{ x -2 }{ 5 }\\\\ 5x+10=2x-4\\\\ 3x=-14\\\\ x= -4\frac{2 }{ 3 } \]

OpenStudy (anonymous):

THANK YOU!!!!!